At Ortman Corporation’ s factory has set the work standards as following. In the factory the labor hours should be 1,050 hours each month to produce 1,500 units of sets of covers at a total direct labor cost of $17,850. The variable overhead is applied at $1 per direct labor hour. The Direct Material standard is 7.9 liters per unit for a total cost of $11,850.

Actual Output 1600 units
Raw Materials Used in Production 12490 liters
Actual Direct labor hours 1090 hours
Purchases of Raw Materials 14000 liters
Actual Price of Raw Materials Purchased $       8.10 per liter
Actual Direct Labor Rate $    16.10 per hour
Actual Variable Overhead Variable Rate $       0.90 per hour

a. The materials quantity variance for May is:

b. The materials price variance for May is:

c. The labor efficiency variance for May is:

d. The labor rate variance for May is:

e. The variable overhead efficiency variance for May is:

f. The variable overhead rate variance for May is:
Answer #1

Solution 1:

Standard quantity of material for actual production = 1600*7.9 = 12640 liters

Actual quantity of material purchased = 14000 liters

Actual quantity of material used = 12490 liters

Standard price of material = $11,850 / (1500*7.9) = $1 per liter

Actual price of material = $8.10 per liter

Material price variance = (SP - AP) * AQ Purchased = ($1 - $8.10) * 14000 = $99,400 U

Material quantity variance = (SQ - AQ) * SR = (12640 - 12490) * $1 = $150 F

Solution 2:

Standard hours per unit = 1050 / 1500 = 0.70 hours per unit

Standard hours of direct labor = 1600 * 0.70 = 1120 hours

Standard rate of direct labor = $17,850/1050 = $17 per hour

Actual hours of direct labor = 1090 hours

Actual rate of direct labor = $16.10 per hour

Direct labor rate variance = (SR - AR) * AH = ($17 - $16.10) * 1090 = $981 F

Direct labor efficiency variance = (SH - AH) * SR = (1120- 1090) * $17 = $510 F

Solution 3:

Standard hours of direct labor = 1600 * 0.70 = 1120 hours

Standard rate of variable overhead = $1 per hour

Actual hours of direct labor = 1090 hours

Actual rate of variable overhead = $0.90 per hour

Variable overhead rate variance = (SR - AR) * AH = ($1 - $0.90) * 1090 = $109 F

Variable overhead efficiency variance = (SH - AH) * SR = (1120 - 1090) * $1 = $30 F